### v1

# Emergence of the Primacy Effect in Structured State-Space Models

Emergence of the Primacy Effect in Structured State-Space Models
Takashi Morita Academy of Emerging Sciences, Chubu University
1 Abstract
Human and animal memory for sequentially presented items is well-documented to be more accurate for those at the beginning and end of a sequence, phenomena known as the primacy and recency effects, respectively. By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time. Accordingly, ANNs are expected to show the recency effect but not the primacy effect. Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture, called structured state-space models, exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
2 Introduction
Human and animal memory for sequentially presented items is well-documented to be more accurate for those appearing at the beginning and end of a sequence—phenomena known as the primacy and recency effects, respectively (Ebbinghaus,, 1913; Murdock,, 1962; Glanzer and Cunitz,, 1966). For example, when a sequence of random integers such as $49,75,\dots,5,38$ is presented in that order, the initial ($49,75$) and final ($5,38$) items are more likely to be recalled accurately at the end of the presentation.
By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time (Bengio et al.,, 1994; Jaeger,, 2001; Jaeger and Haas,, 2004). Thus, ANNs are expected to show the recency effect but not the primacy effect.
Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture—called structured state-space models (Gu et al.,, 2020; Gu et al., 2022b, )—exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
The remainder of this paper is organized as follows. The next section first reviews ANN models for time-series processing. After this preliminary discussion, the Methods section details the task and model specifications of the present study, and the subsequent section presents the results. Finally, the Discussions section summarizes the findings and discusses their implications in relation to previous research.
3 Preliminaries on ANNs for Time-Series Processing
3.1 Recurrent Neural Networks
The traditional ANN-based approach to time-series processing relies on recurrent neural networks (RNNs; Elman,, 1990). RNNs sequentially process inputs, updating their internal state (represented by a real-valued vector) at each time step based on both the current input and the previous state. The output is then generated through a feedforward transformation of this internal state at each step. In other words, RNNs map input time series to target output signals through the latent dynamics of their internal state.
Theoretically, RNNs possess Turing-complete computational power, meaning they can simulate an arbitrary computational model in an “idealized” setting (Siegelmann and Sontag,, 1992, 1994, 1995; Siegelmann,, 1996, 1999). This theoretical capability has driven widespread applications of RNNs in domains such as language modeling (Sundermeyer et al.,, 2012; Graves,, 2013), speech synthesis (Kalchbrenner et al.,, 2018), and weather nowcasting (Shi et al.,, 2015).
However, the “idealized” computational properties of RNNs are unattainable in empirical implementations. For instance, when deployed on standard computers, the weight parameters of RNNs can only be represented with finite precision (Chen et al.,, 2018; Weiss et al.,, 2018), preventing them from storing an unbounded number of observations. Additionally, RNNs exhibit memory decay over time, which has led researchers to explore methods for extending their memory retention capabilities (Hochreiter and Schmidhuber,, 1997; Arjovsky et al.,, 2016; Neil et al.,, 2016; Chang et al.,, 2017; Jing et al.,, 2017, 2019).
3.2 Structured State-Space Models
Recently, researchers have found that continuous-time models can achieve more persistent memory retention than discrete-time RNNs (Zhang et al.,, 2018; Voelker et al.,, 2019; Gu et al.,, 2020). The fundamental principle underlying these models is the polynomial approximation of observed signals (called the HiPPO framework; Gu et al.,, 2020). Specifically, given a single-channel input signal $x(\cdot)$, represented as a function of time, its approximation up to a given time point $t$ can be approximated by a linear combination of polynomials:
$x|_{\leq t}(\cdot)\approx\sum_{n=0}^{N-1}h_{n}(t)P_{n}(\cdot)$
where $P_{n}$ denotes the basis polynomial of degree $n$ and $h_{n}(t)$ are the optimal coefficients for the approximation at time $t$111 The optimal coefficients can be determined as $h_{n}(t):=\langle x|_{\leq t},P_{n}\rangle=\int x|_{\leq t}(s)P_{n}(s)d\mu^{(t)}(s)$ when $\{P_{n}\}_{n=0}^{N-1}$ form an orthogonal basis with respect to the time-dependent measure $d\mu^{(t)}(\cdot)$. Then, these coefficients offer a finite- and constant-dimensional representation of the input signal up to time $t$. The framework can be naturally extended to multi-channel signals by performing channel-wise approximations in parallel, yielding $h_{n}^{(m)}(t)$ for each channel $m$.
For the polynomial coefficients to serve as a “memory” of the input signal, their temporal evolution must be trackable in an online manner; that is, the polynomial approximation at time $t$ should not refer back to past values of the input signal, $x|_{\leq t}(s)$ ($0\leq s<t$). Fortunately, for certain families of polynomials, including Legendre, Laguerre, and Fourier basis,222 The Fourier approximation (or transform) is not based on polynomials, but the theory can be generalized to incorporate it by taking the complex-valued basis $z^{n}:=e^{2\pi inx}$ and a measure on the unit circle (Gu et al.,, 2020). the coefficient dynamics can be described by an ordinary differential equation (ODE; Gu et al.,, 2020):
$\frac{d}{dt}\mathbf{h}(t)=A\mathbf{h}(t)+Bx(t)$ (1)
where $\mathbf{h}(t):=(h_{1}(t),\dots,h_{N}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t):=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
In practice, continuous-time recordings of an input signal $x(t)$ are not available; instead, empirical data consist of discrete-time samples at $t=t_{1},\dots,t_{L}$. Consequently, the SSM matrices must also be discretized in order to convert the ODE in Eq. 1 to a discrete recurrent system, analogous to RNNs:
$\mathbf{h}(t_{j})=\bar{A}\mathbf{h}(t_{j-1})+\bar{B}x(t_{j})$ (2)
where $\bar{A}$ and $\bar{B}$ represent the discretized versions of $A$ and $B$, respectively. A commonly used discretization technique is the bilinear method (Tustin,, 1947), which yields:
$\bar{A}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\left(I+\frac{\Delta t}{2}A\right)$ $\bar{B}:=$ $\left(I-\frac{\Delta t}{2}A\right)^{-1}\Delta tB$
where $\Delta t:=t_{j+1}-t_{j}$ ($\forall j=1,\dots,L-1$) defines the time-step size. This time-step parameter is treated as learnable (for each input channel), allowing the model to automatically adjust the time scale of its state-space dynamics to align with that of the input signal.
This flexibility extends the applicability of SSMs to inherently discrete data lacking overt continuous dynamics (e.g., text languages); the model can jointly lean latent representations (or embeddings) of discrete inputs along with their (pseudo-)continuous dynamics via gradient-based optimization. Furthermore, SSMs can be hierarchically stacked to construct deeper and more expressive ANNs, where each layer processes latent signals received from lower layers.
Figure 1: Illustration of the continuous (left; $K(t):=e^{tA}B$) and discretized (right; $\bar{K}_{-j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). The polynomial degree is indicated by the line colors. The kernels were evaluated at $t=-j\Delta t$ for $j=0,\dots,-255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.
Previous studies have identified that the time-step size, $\Delta t$, as a critical factor in determining the success/failure of SSMs. Intuitively, a small $\Delta t$ results in minor state updates, yielding slow dynamics in the state space, $\mathbf{h}(t_{j+1})-\mathbf{h}(t_{j})$; conversely, a large $\Delta t$ induces rapid state transitions (Gu et al.,, 2021; Gu and Dao,, 2024). As a consequence, $\Delta t$ governs the memory decay properties of the SSM; although the models assume an exponentially decaying measure in continuous time (particularly when employing Legendre/Laguerre polynomials; Gu et al.,, 2023), choosing small $\Delta t$ values can allocate relatively large weights to input samples from distant time steps, thereby compromising single-step discriminability, which is better preserved with larger $\Delta t$ (see left panels in Fig. 1).
It is also important to remember that discretization methods are generally designed under the assumption of sufficiently small $\Delta t$, for effectively approximating the limit $\Delta t\to 0$. Consequently, setting $\Delta t$ too large introduces discrepancies between the continuous and discretized dynamics (as illustrated by the contrast between the left vs. right panels in Fig. 1).
3.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at the current time $t$ while maintaining full access to the most recent $M$ observations, $x_{t-M-1},\dots,x_{t}$. At each time step, they “forget” or discard the oldest observation ($x_{t-M}$) as new data enter the storage buffer.
By definition, finite-queue/Markov models do not exhibit the primacy effect, as the earliest observations are inevitably lost once they exceed the finite storage window. Instead, these models align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). Given this fundamental limitation, finite-queue/Markov models fall outside the scope of the present study.
[…]
Figure 2: Schematic illustration of the binary memory verification task.
4 Methods
4.1 Task
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).444 The most well-adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in a theoretically more demanding task requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
The task hyperparameters were manually adjusted to prevent the models from achieving perfect accuracy. Specifically, the input length was set to $L\in\{64,128,256\}$, and the vocabulary size was fixed at $K:=4096$. Each model underwent ten independent training runs with different random seeds. For evaluation, 1024 sets of integers were held out as test data, ensuring that these integer combinations never appeared as study items in the training set, regardless of their order.
To build test sequences, the held-out study items were randomly ordered, and queries were first generated by first shuffling and then cyclically shifting them (e.g., $(2,8,11,29)\mapsto\{(2,8,11,29),(8,11,29,2),(11,29,2,8),(29,2,8,11)\}$). This design ensured that each study item was queried in all $L$ possible positions. Finally, either the even- or odd-indexed query positions were replaced with random distractors, resulting in a total of $1024\times L\times 2$ test sequences per trial.
4.2 Models
The models used for the binary memory verification task comprised three layers, as illustrated in Fig. 2. First, the input integers were embedded into 256-dimensional real-valued vectors. These embeddings were shared between study items and query tokens. The resulting sequence of vectors was then processed by the SSM/RNN, whose outputs were linearly projected onto binary logits to determine whether each query token was present in the study items.
This study primarily examined the single-layer S4 model as the goldstandard implementation of the SSM (Gu et al., 2022b, ).555 Recent studies have shown that the state matrix ($A$) of S4 can be simplified into a purely diagonal form without compromising performance (S4D; Gu et al., 2022a, ). By contrast, the original S4 model introduced an additional low-rank component to the diagonal structure (referred to as the Diagonal Plus Low Rank form, or DPLR) to ensure a mathematically well-founded state matrix. Notably, the diagonal variant exhibited a qualitatively similar primacy effect to the DPLR model. Due to the page limitations, results for the diagonal model are omitted from this paper, and all reported findings are based on the DPLR model. The model encoded the channel-wise dynamics of the input embeddings in a complex-valued space, with its outputs subsequently projected back into the real domain by discarding imaginary components. The state and input matrices ($A$ and $B$ in Eq. 1) were initialized to approximate each channel’s trajectory using Legendre/Laguerre polynomials of degrees 0–63 (HiPPO-LegS/LagT) or a Fourier basis $\{s_{0},c_{0},\dots,s_{31},c_{31}\}$, where $s_{n}(t):=\sqrt{2}\sin(2\pi nt)$ and $c_{n}(t):=\sqrt{2}\cos(2\pi nt)$ (HiPPO-Fout, Fourier Recurrent Unit; Zhang et al.,, 2018; Gu et al.,, 2020, 2023). The matrices were discretized by the bilinear method (Tustin,, 1947).
For comparison, a single-layer long short-term memory (LSTM) network was also evaluated (Hochreiter and Schmidhuber,, 1997). LSTM has been the goldstandard RNN architecture for various time-series processing tasks, including language modeling (Sundermeyer et al.,, 2012; Graves,, 2013). The dimensionality of both hidden and cell states of was set to 256.
The models were trained for 300,000 iterations using the Adam optimizer with parameters $(\beta_{0},\beta_{1}):=(0.9,0.99)$ (Kingma and Ba,, 2015). Batch size was set to 512. The learning rate was linearly increased from 0.0 to 0.001 over the first 1,000 iterations (warmups) and subsequently decayed according to the cosine annealing schedule Loshchilov and Hutter, (2017). To prevent gradient explosion, the gradient norm was clipped at 1.0. The Python code for the experiments is available at https://github.com/An0nym0usAuth0r/NeuralPrimacyEffect.git.
5 Results
5.1 Emergence of the Primacy Effect
[…]
Figure 3: Accuracy of the binary memory verification task. Each cell in the heatmap represents the accuracy of study items that were first observed at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the S4 model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ ($0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
The binary memory verification performance of the S4 model was highest for study items presented at the beginning of the sequence, demonstrating a clear primacy effect (Fig. 33). Additionally, the model maintained high accuracy across different query timings (as indicated by the bright colors in the top rows of the heatmaps), provided that the sequence length did not exceed its capacity (see the accuracy decline in Fig. 3, where $L=256$). In other words, memory for the initial study items exhibited minimal decay over time.
By contrast, the LSTM did not display this primacy effect; its accuracy was uniform across both the memorization and verification phases (Fig. 3).
Interestingly, the S4’s accuracy for the most recently presented study items was lowest when they were queried immediately after their initial presentation in the memorization phase (indicated by the dark colors in the bottom-left region of the heatmaps). This suggests a temporal delay between the encoding of study items and their effective retrieval.
These findings held true regardless of whether the state and input matrices of the SSM were optimized for the task (Fig. 3) or remained fixed at their initial values (Fig. 33). Moreover, the results remained consistent across different polynomial bases underlying the state and input matrices, including Laguerre (Fig. 3), Fourier (3), and Legendre (all other panels).
Figure 4: Top row: Optimization trajectories of the discretization step size $\Delta t$ in the S4 model with frozen Legendre state and input matrices. Each heatmap column represents the distribution of $\Delta t$ across 256 latent channels $\times$ 10 training runs, sorted in the ascending order. Bottom row: Histograms displaying the initial (blue) and final (orange) values of $\Delta t$. The first to third panel columns from the left show the results for three different ranges of log-uniformly random initializations, whereas the fourth and fifth columns tested shorter and longer study items, respectively.
5.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ is a key parameter governing the memory capacity of the SSM, although the primacy effect emerged across different initial values (Fig. 3). To further investigate its role, the learning trajectories of $\Delta t$ over training iterations were analyzed. The results indicated that as training progressed, a specific range of step sizes became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig 4).
Additionally, the peak value of $\Delta t$ was found to depend on the number of the study items $L$; longer study sequences led the model to favor smaller $\Delta t$ values (compare the leftmost panel with the two rightmost panels).
6 Discussions
6.1 Summary of the Findings
The present study demonstrated that the SSM exhibits the primacy effect in memorization. When performing the binary memory verification task, which parallels paradigms used to investigate memory capacities in humans and animals (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985), the model showed the highest accuracy for study items presented at the beginning of the sequence. Moreover, memorized information was not retrievable immediately after the presentation of the study items. These findings are novel and counterintuitive, as they challenge the theoretical formulation of the SSM, which assumes an exponentially decaying measure (for Legendre and Laguerre bases; Fig 1; Gu et al.,, 2023).
As noted in the Preliminaries, the SSM was designed to achieve a longer-lasting memory than classical RNNs (Gu et al.,, 2020). Prior research on RNNs and SSMs has focused on their ability to preserve input data against temporal decay. A typical benchmark for these inquiries is the delayed reconstruction task (a.k.a. the copying memory task; Arjovsky et al.,, 2016). In this task, models are provided with a short sequence of integers—usually ten repeatable random samples ranging between 0 and 7—followed by null inputs (e.g., zero vectors), and are then asked to reconstruct the initially observed sequence in the original order. Model performance is evaluated based on the length of the null-valued interval between the initial observations and their reconstructions.
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored.666 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
6.2 Comparison with Data/Task-Oriented Primacy Effects
Several prior studies have reported a primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) examined positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.777 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: ”) consistently corresponded to their order of presentation. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels.
Xiao et al., (2024) found that Transformer-based LLMs allocated disproportionate attention to initial tokens, regardless of their informational salience in the text (a phenomenon they termed attention sinks). Furthermore, retaining these initial tokens even after they fall outside the prespecified input window enhanced model performance.
Since the Transformer architecture is inherently position-agnostic—lacking an intrinsic ordering mechanism apart from external positional encodings (Vaswani et al.,, 2017)—the primacy effects observed in these studies must stem from the structure of the training data or task design. Wang et al., argued that LLMs inherit cognitive biases from human-generated linguistic data. Xiao et al., suggested that the nature of the language modeling task itself encourages prioritization of initial tokens, as they are repeatedly used as inputs for autoregressive predictions, reinforcing attention allocation to them.
By contrast, the findings of the present study are based on a synthetic memorization task designed to closely resemble controlled psychological experiments (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). Therefore, the observed primacy effect is attributed to the intrinsic properties of the SSM per se, rather than to biases introduced by data or task design.
6.3 Implications for Computational Neuroscience
Classical theories on the primacy effect in human and animal memory hypothesize a dual-storage mechanism, attributing the primacy and recency effects to long- and short-term memory systems, respectively (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). According to these theories, initially presented study items are transferred from short-term memory—where all incoming observations are temporally buffered—into long-term memory, where they are preserved against decay, thereby explaining for the primacy effect. By contrast, the recency effect arises from the residual information retained in short-term memory, which has a limited capacity and can store the most recently presented items.
By contrast, the findings of the present study may contribute to the development of a novel and more neurobiologically grounded model of the primacy effect. Voelker and Eliasmith, (2018) and Voelker et al., (2019) established a connection between the SSM and spiking neural networks, demonstrating that the dynamics of the state vector in the SSM (Eq. 1) can be implemented using a population of spiking neurons. Given that spiking neural networks remain the most plausible model of biological neuronal activity to date, a spiking version of the SSM studied here could serve as a new neurobiological model of the primacy effect.
Furthermore, Voelker and Eliasmith, (2018) also showed that the spiking representation of an SSM, when tuned to implement some particular function known as the delay line, replicates the behavior of biological time cells,888 Time cells fire at specific moments following onset cues/events and are believed to encode temporal information (MacDonald et al.,, 2011). which have been observed in several brain regions, including the hippocampus (MacDonald et al.,, 2011; Eichenbaum,, 2014), striatum (Mello et al.,, 2015), and cortex (Luczak et al.,, 2015). In such a way, comparing a spiking implementation of the SSM that exhibits the primacy effect with brain recordings from psychological experiments could provide novel computational insights into the neuronal mechanisms underlying this cognitive bias.
Of course, the present study has certain limitations as a biological simulation. Most notably, the model here was explicitly optimized for the memorization task from scratch, whereas the primacy effect observed in humans and animals is presumably an outcome of their natural cognitive development. Thus, fundamentally different mechanisms may underlie the similar memorization biases exhibited by the SSM and biological organisms. Future studies may address this discrepancy by training the model in a more biologically naturalistic setting. However, as discussed in the previous section, training an ANN on empirical data, such as natural language, introduces the risk of conflating human-originated cognitive biases—transferred through the data—with the model’s intrinsic properties.
The model’s capacity also far exceeded that of biological memory systems. The binary memory verification task employed in this study involved extremely long lists of study items ($L\in\{64,128,256\}$) as well as a huge vocabulary ($K=4096$), which is unlikely to be replicated in human or animal experiments. This level of task difficulty was necessary to clearly demonstrate the primacy effect in the SSM. However, future research should explore downscaling the model toward more biologically plausible conditions.
Finally, the present study does not offer an explanation for various well-documented properties of the psychological primacy effect. For instance, the retention of initial study items is thought to be achieved, or at least facilitated, by rehearsal—repetitive mental or verbal recitation of observed items before requested recall/verification (Rundus,, 1971; Marshall and Werder,, 1972; Glenberg et al.,, 1980)—for which the models studied here lack an analogous mechanism. In addition, the models operated on a unitless discrete time scale, to which the continuous-time representation of the SSM was aligned. As a result, they are unsuitable for investigating physical-time effects, such as the impact of retention intervals between study-item presentation and verification (Cornell and Bergstrom,, 1983; Wright et al.,, 1985; Neath,, 1993) or variations in inter-item intervals (Neath and Knoedler,, 1994). Future studies may address this latter issue by employing more biologically realistic neural network models; as noted above, the SSM is closely related to spiking neural networks, which are capable of simulating physical-time dynamics and incorporating biologically inspired connectivity structures (Voelker and Eliasmith,, 2018).
7 Acknowledgments
This study was supported by JST ACT-X (JPMJAX21AN) and Core Research for Evolutional Science and Technology (JPMJCR22P5); JSPS Grant-in-Aid for Early-Career Scientists (JP21K17805) and for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the ACCMS, Kyoto University, regarding the use of their supercomputer system.